# Multiple Choice Learning of Low Rank Adapters for Language Modeling

### v1

@@ 5.3 Image description with reading comprehension
5.3.1 Quality and Diversity Evaluation
Similarly to Figure 2, Table 2 confirms the trends observed in the audio captioning task. At an equal number of forward passes, LoRA-MCL outperforms LoRA-MLE, even using DBS decoding with a $\lambda$ diversity parameter specifically optimized for the task (SPIDEr of $0.955$ against $0.915$). However, we noticed that the DBS decoding of LoRA-MLE tends to generate more diverse outputs than the greedy decoding of LoRA-MCL (mBLEU of $0.416$ against $0.520$). Interestingly, increasing the number of beams in each group decreases diversity and does not improve the performance of LoRA-MLE, while with LoRA-MCL, increasing the beam size (from 1 to 2) not only slightly increases the performance but also improves the diversity of the generations, indicating promising further inference-time scaling. Additional results are provided in Appendix (Table 9).
To highlight the behavior of LoRA-MCL in a realistic case where one can control the modes of the data-generating process, we simulated an artificial bi-modal distribution of the dataset, similarly to the setup of the toy experiment of Section 4.3. We did so by translating half of the captions of the data from English to French (using T5-small T5 ), while keeping the prompts in English.
We trained and evaluated a 2-hypothesis model in the LoRA-MCL setup, as well as a baseline trained in LoRA-MLE. We observed a clear specialization of each hypothesis towards a given language (i.e., one hypothesis learned French and the other English): at test-time, the winning head is the first one in $\sim 89\%$ of the French captions and the second one in $\sim 97\%$ of the English captions. Table 3 shows performance and diversity scores of both LoRA-MLE and LoRA-MCL on our synthetic test set. LoRA-MCL is decoded greedily while LoRA-MLE uses DBS with the $\lambda$ value maximizing its performance (0.8). Although the overall performances are similar, LoRA-MCL generates significantly more diverse captions (mBLEU-4 of 0.027 and 0.029 against 0.138 and 0.126 on the French and English subparts, respectively), and outperforms LoRA-MLE in the French part of the test set (SPIDEr of 0.464 against 0.411).
These results relate to the synthetic data experiments of Sec. 4.2: in the presence of two modes, LoRA-MLE learns an average of them. Here, we believe this is more shifted toward English as the original LLaVA model performs better in English than in French. In contrast, LoRA-MCL is expected to learn the two modes separately. We believe this is what explains the gain in performance in French, at the cost of a slight reduction in English performance.
Fig. 3 qualitatively illustrates the behavior of the models: LoRA-MLE learns a weighted average of the two modes, shifted towards English, and it is sometimes not able to output French captions within the 2 candidates as in the book example. We found LoRA-MLE to be particularly prone to errors when outputting French sentences: for instance, in the second generation of examples with an image of a bottle, LoRA-MLE enters a repetition loop. On the other hand, LoRA-MCL is less affected by those artifacts on French sentences, and successfully captures the two modes of the distribution, benefiting from the specialization of each hypothesis.
Table 3: SPIDEr ()($\uparrow$)( ↑ ) & mBLEU-4 ()($\downarrow$)( ↓ ) on different parts of synthetic test set.
@@ 5.3.2 Observing hypothesis specialization in bilingual image description
Figure 3: Observing specialization in bilingual image description. Quantitative (Left) and Qualitative (Right) analysis for LoRA-MLE and LoRA-MCL in the setup of Section 5.3.2.
@@ 6 Related Work
Multiple Choice Learning to predict diverse and plausible outputs. Multiple Choice Learning guzman2012multiple ; lee2016stochastic is a training paradigm that minimizes the oracle or winner-takes-all loss across a set of models, encouraging specialization rupprecht2017learning ; letzelter2024winner . While it has been noted that the collapse issue needs to be addressed with care rupprecht2017learning ; amcl , MCL has demonstrated broad applicability across machine learning tasks lee2017confident ; seo2020trajectory ; garcia2021distillation , typically using a shared backbone and multiple projection heads. However, using multiple heads may be impractical in large language models. Mixture-of-Experts jacobs1991adaptive ; shazeer2017outrageously can be an effective strategy to manage computational costs in this case. However, as primarily thought for model scalability and not diversity, only a small subset of experts are active at each forward pass via routing, making them incompatible with MCL, while still suffering from knowledge redundancy in the experts jiang2024mixtral . To the best of our knowledge, we are the first to adapt MCL to the setup of language modeling with next-token prediction, specifically, using internal models instantiated via multiple LoRA hu2022lora .
@@
Generating Multiple Outputs with Language Models. Language models are commonly trained via next-token prediction, framed as Maximum Likelihood Estimation. This is arguably the most popular method for training large-scale language models shlegerislanguage ; radford2019language ; touvron2023llama , including those that take audio or images as input chu2024qwen2 ; liu2023visual , with much of its success attributed to tokenization makkuva2024attention ; rajaraman2024analysis . Generating diverse and plausible sequences at inference remains challenging: $(i)$ Sampling methods holtzmancurious ; fan2018hierarchical ; meister2023locally may be unreliable depending on the chosen parameters; $(ii)$ Exact MAP decoding is intractable due to the exponential search space eikema2020map ; $(iii)$ Strategies like Beam Search often yield repetitive or overly coherent outputs fan2018hierarchical ; holtzmancurious ; keskar2019ctrl . Diverse Beam Search mei2022diverse or contrastive decoding sucontrastive inject diversity through test-time parameters (e.g., penalty $\lambda$), but in contrast, our method infers diversity from the input’s inherent ambiguity.
@@
Diversity in Audio and Visual Captioning. Audio drossos2017automated ; mei2021audio ; mei2022automated and image captioning aneja2018convolutional ; herdade2019image ; hossain2019comprehensive have traditionally relied on dataset- and task-specific models trained with Maximum Likelihood Estimation (MLE). A key challenge identified in both domains is the lack of diversity in generated captions, prompting the development of other training objectives specifically designed to address this issue mei2022diverse ; xu2022diversity ; xu2024towards ; zhang2024generating ; wang2019describing ; wang2020diversity ; mahajan2020diverse . These methods often require architectural modifications and are typically designed for models trained from scratch. However, recent advances have shifted the field toward general-purpose multimodal large language models chu2024qwen2 ; liu2023visual , which offer broader capabilities but pose new challenges to address the diversity-quality trade-off. In this work, we demonstrate that LoRA-MCL can be effectively applied at the fine-tuning stage of such models.
@@ 7 Conclusion
This paper proposes a new paradigm for training language models to generate diverse plausible predictions. This is done with LoRA-MCL, combining the Multiple Choice Learning paradigm and Low-rank Adaptation. We provide a theoretical analysis showing that when the target sequence is written as a mixture, LoRA-MCL is expected to capture the modes of the data distribution. We validate our analysis with experiments on Markov chains, and on the real-world audio and advanced image captioning tasks. Future work includes a careful interpretation of the concepts learned by each model as a function of the ambiguity of the input, e.g., for uncertainty estimation as an alternative to wang2023lora .
Limitations. In LoRA-MCL, setting $\varepsilon$ too high ensures that all hypotheses receive gradients, but it can also overly homogenize the models, thereby reducing diversity. While MCL performs well across a broad range of $\varepsilon$ values, identifying the optimal setting can be challenging. To do so, we plan to incorporate recent annealing MCL techniques, e.g., amcl , which employ a scheduling mechanism for $\varepsilon$ and may help improve the quality of the solution. Note that recent LoRA variants, such as hayou2024lora+ may also be an orthogonal axis of improvement for our method.
Acknowledgments
We thank Mickael Chen for his valuable advice, as well as Gilles Puy for his help in the implementation of the parallelization over the hypotheses. This work was funded by the French Association for Technological Research (ANRT CIFRE contract 2022-1854), and granted access to the HPC resources of IDRIS (allocation 2024-AD011014345) by GENCI.
@@ Appendix A Notations and setup
In the following, $x\triangleq(x_{t})_{t=1}^{T}\in\mathcal{V}^{T}$ be a sequence of $T$ tokens belonging to a finite vocabulary $\mathcal{V}=\{1,\dots,|\mathcal{V}|\}$, and $c\triangleq(c_{t})_{t=1}^{\tau}\in(\mathbb{R}^{d})^{\tau}$ be a sequence of $\tau$ context of embeddings of dimension $d$. In the following $\mathcal{X}=\mathcal{V}^{T}$ and $\mathcal{C}=(\mathbb{R}^{d})^{\tau}$.
@@
In the following, we assume that the model $p_{\theta}$ is perfectly expressive. Formally, let $\mathcal{F}_{\theta}\triangleq\{p_{\theta}:\mathcal{C}\rightarrow\mathcal{P}(\mathcal{X})\}$ be the family of conditional distributions realized by the model, where $\mathcal{P}(\mathcal{X})$ is the set of probability distributions on $\mathcal{X}$. We assume the family $\mathcal{F}_{\theta}$ is perfectly expressive, that is $\mathcal{F}_{\theta}=\{p^{\star}:\mathcal{C}\rightarrow\mathcal{P}(\mathcal{X})\}$.
First, note that we have the Proposition 2, a well-known result that is due to the Gibbs inequality (e.g., [42]), for which we provide a proof for completeness.
Proposition 2 ([42]).
@@
where $\mathcal{H}(x\,|\,c)\triangleq-\mathbb{E}_{c,x}[\mathrm{log}\;p(x\,|\,c)]$ is the entropy of $p(\cdot\,|\,c)$.
@@
$-\int_{\mathscr{S}(p)}\mathrm{log}\left[\frac{p_{\theta}(x\,|\,c)}{p(x,c)}\right]p(x,c)\;\mathrm{d}x\mathrm{d}c\geq-\mathrm{log}\left(\int_{\mathscr{S}(p)}\frac{p_{\theta}(x\,|\,c)}{p(x,c)}p(x,c)\;\mathrm{d}x\mathrm{d}c\right)\geq 0\;.$ (14)
@@
$-\int_{\mathcal{X}\times\mathcal{C}}p(x,c)\;\mathrm{log}\;p_{\theta}(x\,|\,c)\;\mathrm{d}x\mathrm{d}c\geq-\int_{\mathcal{X}\times\mathcal{C}}p(x,c)\;\mathrm{log}\;p(x\,|\,c)\;\mathrm{d}x\mathrm{d}c=\mathcal{H}(x\,|\,c)\;,$
@@
$-\mathbb{E}_{\alpha}[\mathrm{log}\,\beta(x)]=\mathrm{KL}(\alpha\,\|\,\beta)+\mathcal{H}(\alpha)\;,$ (15)
@@ Appendix B Proof of Proposition 1
$D(\theta,q)\triangleq\int_{\mathcal{X}}\sum_{k=1}^{K}q(k\,|\,x)\mathrm{log}\,p(x,z_{k};\theta)\;\mathrm{d}p(x),\leavevmode\nobreak\ D(\theta)\triangleq\int_{\mathcal{X}}\max_{k=1,\dots,K}\;\mathrm{log}\,p(x,z_{k};\theta)\;\mathrm{d}p(x)\;,$ (18)
@@
Indeed, the expectation step comes down to computing $q(\theta^{(t)})$. For for the Maximization step, we have: $D\left(\theta^{(t+1)},q(\theta^{(t)})\right)\geq D\left(\theta^{(t)},q(\theta^{(t)})\right)$, by definition. At the next expectation step, we have: $D\left(\theta^{(t+1)},q(\theta^{(t+1)})\right)\geq D\left(\theta^{(t+1)},q(\theta^{(t)})\right)$, because $q(\theta^{(t+1)})$ computes the best assignment given the parameters $\theta^{(t+1)}$. This shows that $D(\theta^{(t+1})\geq D(\theta^{(t)})$.
First note that the main difference compared to the vanilla form of the (hard) EM algorithm, is that the goal is to fit here a conditional distribution $p(x\,|\,c)$ given pair $(c,x)\sim p(c,x)$. Furthermore, step 2 performs a gradient step (of the neural network weights) instead of a full maximization.
@@
$\mathcal{L}^{\mathrm{WTA}}(\theta)=-\int_{\mathcal{X}\times\mathcal{C}}\max_{k=1,\dots,K}\mathrm{log}\;p(x\,|\,c;\theta_{k})\mathrm{d}p(c,x)\;,$ (19)
which is expected to decrease at each training iteration. Note that because the loss is bounded from below (by 00), the sequence of real numbers $\{\mathcal{L}^{\mathrm{WTA}}(\theta^{(t)})\}_{t\geq 0}$ is therefore expected to converge.
To conclude, we can view $(\theta_{1},\dots,\theta_{K})$ as the parameters involved in the estimation of the modes of the conditional distribution with $p(x\,|\,c)=\sum_{k=1}^{K}p(x\,|\,c;\theta_{k})p(\theta_{k}\,|\,c)$. Note that the current form of the algorithm does not estimate the weight of each mode $p(\theta_{k}\,|\,c)$, further work could include incorporating scoring heads to estimate $p(\theta_{k}\,|\,c)$ each $k$ as in [35]. ∎
We then expect to be able to recover the distributions (with one permutation) $\{p(x\,|\,c,z_{k})\}$ from estimated $\{p(x\,|\,c;\theta_{k})\}$, assuming identifiability of the data generating mixture, which we expect to be made easier if the components are enough separated (See e.g., [59, Par. 2.5] or [47, Sec. 2.2]).
Proof of (ii) Let us assume that (with one permutation) $p(x\,|\,z_{k},c)=p(x\,|\,c;\theta_{k})$ for $k\in\{1,\dots,K\}$. This is possible thanks to Assumption 1. Let us show that (16).
Let us define:
@@
Now let us leverage Assumption 1 to choose $\tilde{\theta}_{k}$ such that $p(x\,|\,c,\tilde{\theta}_{k})=p(x\,|\,c,z_{k})$ for each $k\in\{1,\dots,K\}$. In this case, $\mathrm{max}_{k}\,p(x\,|\,c,\tilde{\theta}_{k})\geq p(x\,|\,c,z)$ for each $z\in\{z_{1},\dots,z_{K}\}$, and $-\mathbb{E}_{z}\mathbb{E}_{x\,|\,z}\left[\mathrm{log}\frac{\mathrm{max}_{k}\,p(x\,|\,c,\tilde{\theta}_{k})}{p(x\,|\,c,z)}\right]\leq 0$. Then,

### v2

@@ 5.3 Image description with reading comprehension
Table 2: SPIDEr $(\uparrow)$ & mBLEU-4 $(\downarrow)$ on different parts of synthetic test set.
@@
Figure 4: Observing specialization in bilingual image description. Quantitative (Left) and Qualitative (Right) analysis for LoRA-MLE and LoRA-MCL in the setup of Sec. 6.
Table 3: Quality and Diversity Evaluation on TextCaps (3 candidates). Best in bold; second-best underlined. Higher is better $(\uparrow)$ except mBLEU-4 $(\downarrow)$.
Training Decoding Beam $\mathrm{mBLEU}_{4}$ $\mathrm{CIDEr}_{D}$ $\mathrm{SPICE}$ $\mathrm{SPIDEr}$
LoRA-MLE BS 3 0.688 1.517 0.244 0.873
LoRA-MLE BS 6 0.786 1.557 0.246 0.895
LoRA-MLE DBS ($\lambda=0.8$) 3 0.437 1.590 0.251 0.909
LoRA-MLE DBS ($\lambda=1.0$) 3 0.416 1.586 0.250 0.906
LoRA-MLE DBS ($\lambda=0.8$) 6 0.671 1.573 0.251 0.903
LoRA-MLE DBS ($\lambda=0.8$) 3 0.531 1.589 0.255 0.912
LoRA-MLE DBS ($\lambda=1.0$) 3 0.425 1.601 0.252 0.915
LoRA-MoE DBS ($\lambda=0.8$) 3 0.441 1.616 0.254 0.924
LoRA-MoE DBS ($\lambda=1.0$) 3 0.421 1.622 0.253 0.926
LoRA-MoE DBS ($\lambda=0.8$) 6 0.678 1.608 0.255 0.922
LoRA-MCL BS 1 0.520 1.674 0.255 0.955
LoRA-MCL BS 2 0.490 1.627 0.258 0.932
Figure 5: Quality vs. Diversity in Machine Translation.
5.3.1 Quality and Diversity Evaluation
Evaluation in image captioning Table 3 confirms the trends observed in the audio captioning task. At an equal number of forward passes, LoRA-MCL outperforms LoRA-MLE and LoRA-MoE, even using DBS with a $\lambda$ diversity parameter specifically optimized for the task (SPIDEr of $0.955$ vs. $0.926$). Consistently with audio captioning, increasing the number of beams in each group can decrease diversity and does not improve the performance of LoRA-MLE and LoRA-MoE. However, we noticed that the DBS with LoRA-MLE tends to generate more diverse outputs than the greedy decoding of LoRA-MCL (mBLEU of $0.416$ vs. $0.520$). Combining DBS with LoRA-MCL, inspired by Guzman-Rivera et al. (2014), could help address this issue. Additional comparisons are provided in Apx. E.7.2.
Figure 6: PCA of hypothesis candidate embeddings. PC1 and PC2 correspond to the first two principal components, explaining $31.81\%$ and $15.22\%$ of the variance respectively.
To highlight the behavior of LoRA-MCL in a realistic case where one can control the modes of the data-generating process, we simulated an artificial bi-modal distribution of the dataset, similarly to the setup of the toy experiment of Sec. 4.3. We did so by translating half of the captions of the data from English to French (using T5-small (Raffel et al., 2020)), while keeping the prompts in English.
We trained a two-hypothesis LoRA-MCL model and LoRA-MLE baseline. We observed a specialization of each hypothesis towards a given language (one hypothesis learned French and the other English): at test-time, the winning head is the first one in $\sim 89\%$ of the French captions and the second one in $\sim 97\%$ of the English captions. Table 4 reports quality/diversity on the synthetic test set. LoRA-MCL uses greedy decoding, LoRA-MLE DBS with $\lambda=0.8$ (maximizing its performance). Overall performance is similar, but LoRA-MCL is notably more diverse (mBLEU-4: FR $0.027$ vs $0.138$, EN $0.029$ vs $0.126$) and outperforms on French (SPIDEr $0.464$ vs $0.411$), with a slight reduction in English performance. Consistently with Sec. 4.2, LoRA-MLE learns an average of the two modes (likely biased to English from LLaVA pretraining), whereas LoRA-MCL separates them. Additional experimental evidence is provided in Apx. E.7.4.
Fig. 4 illustrates the behavior of the models: LoRA-MLE learns a weighted average of the two modes, shifted towards English, and sometimes fails to output French captions within the two candidates, as in the book example. We found LoRA-MLE to be prone to errors when outputting French sentences: in the generation examples with an image of a bottle, LoRA-MLE enters a repetition loop. On the other hand, LoRA-MCL is less affected by those artifacts on French sentences, and captures the two modes of the distribution, benefiting from hypothesis specialization. Fig. 6 presents a PCA of the captions generated by each head, embedded using a Sentence-BERT model (StyleDistance from Patel et al. (2025)). The resulting frontiers are clearly visible, demonstrating a strong specialization of each head. See Apx. E.7.4 for further analysis.
5.4 Diverse Machine Translation
We evaluate LoRA-MCL for zero-shot machine translation with LLMs. Following Xu et al. (2024a), we use a two-stage paradigm: (1) full-parameter fine-tuning on a monolingual corpus, and (2) LoRA fine-tuning on parallel data. We build on ALMA-7B, stage-1 fine-tuned, which we fine-tune on parallel data from Xu et al. (2024a) (see Apx. E.8), comprising WMT’17–20 test sets and Flores-200 dev/test sets, restricted to English–German ($\sim$14k pairs). Evaluation uses the newstest2014 subset from Ott et al. (2018), containing 500 English sentences with ten German references each.
Fig. 5 displays the results, with a legend mirroring the one in Fig. 3. LoRA-MCL (circles) is trained with $K=3$ and $\varepsilon=0.05$, and each model generates $3$ sequences per input. We evaluate LoRA-MLE with ranks $r=16$ (diamond) and $r=48$ (square), using BS (blue) with widths $3$, 66, and $9$, and DBS (green) with $\lambda=0.8$. We follow the quality–diversity evaluation protocol of Shen et al. (2019). Scores are reported using both Leave-One-Out BLEU (Loo-BLEU) and Pairwise-BLEU (Shen et al., 2019). The results show that LoRA-MCL achieves a balance between quality and diversity, confirming the effectiveness of the method.
@@ 6 Related Work
MCL to predict diverse and plausible outputs. MCL (Guzman-Rivera et al., 2012; Lee et al., 2016) is a training paradigm that minimizes the WTA loss across a set of models, encouraging specialization (Rupprecht et al., 2017; Letzelter et al., 2024). While the collapse issue needs to be addressed (Rupprecht et al., 2017; Perera et al., 2024), MCL has demonstrated broad applicability across tasks (Lee et al., 2017; Seo et al., 2020; Garcia et al., 2021), typically using a shared backbone and multiple heads. However, using multiple heads may be impractical in LLMs. Mixture-of-Experts (MoE) (Jacobs et al., 1991; Shazeer et al., 2017) offers an alternative for managing computational costs, since only a subset of experts is active at each forward pass. In LLMs, however, MoE has primarily been used to improve scalability rather than to encourage diversity, and suffers from redundancy among experts (Jiang et al., 2024). While MoE can be adapted to the LoRA setting (Wu et al., 2024; Li et al., 2024), there is no clear consensus on the degree of specialization achieved. To our knowledge, this work is the first to adapt MCL to next-token language modeling using multiple LoRA modules.
@@
Generating Multiple Outputs with LMs. Language models are commonly trained via next-token prediction, framed as Maximum Likelihood Estimation. This is arguably the most popular method for training large-scale language models (Shlegeris et al., 2022; Radford et al., 2019; Touvron et al., 2023), including those that take audio or images as input (Chu et al., 2024; Liu et al., 2023), with much of its success attributed to tokenization (Makkuva et al., 2024; Rajaraman et al., 2024). Generating diverse and plausible sequences at inference remains challenging: $(i)$ Sampling methods (Holtzman et al., 2019; Fan et al., 2018; Meister et al., 2023) may be unreliable depending on the chosen parameters; $(ii)$ Exact MAP decoding is intractable due to the exponential search space (Eikema and Aziz, 2020); $(iii)$ Strategies like Beam Search often yield repetitive or overly coherent outputs (Fan et al., 2018; Holtzman et al., 2019; Keskar et al., 2019). Diverse Beam Search (Mei et al., 2022b) and Test-time Augmentation (Wang et al., 2019; Kim et al., 2022; Kaya et al., 2025) inject diversity through test-time parameters (e.g., penalty $\lambda$), but in contrast, LoRA-MCL infers diversity from input’s ambiguity.
@@
Diversity in Audio and Visual Captioning. Audio (Drossos et al., 2017; Mei et al., 2021, 2022a) and image captioning (Aneja et al., 2018; Herdade et al., 2019; Hossain et al., 2019) have traditionally relied on MLE-trained, task-specific models. A key challenge is the limited diversity of generated captions, leading to training objectives tailored to this issue (Mei et al., 2022b; Xu et al., 2022, 2024b; Zhang et al., 2024; Wang and Chan, 2019; Wang et al., 2020; Mahajan and Roth, 2020). These approaches often require architectural changes and models trained from scratch. With the rise of general-purpose multimodal LLMs (Chu et al., 2024; Liu et al., 2023), addressing the diversity–quality trade-off remains critical. We show that LoRA-MCL effectively tackles this at the fine-tuning stage.
@@ 7 Conclusion
LoRA-MCL combines MCL with LoRA to train language models for diverse, plausible predictions. We show that when the target sequence is a mixture, LoRA-MCL can capture the modes of the data distribution. We validate this on Markov chains as well as the tasks of audio, image captioning, and machine translation. Future work will focus on further interpreting the concepts learned by each model, and investiging recent recent LoRA variants such as Hayou et al. (2024); Zhang et al. (2025).
Limitations. In the relaxed variant, setting $\varepsilon$ too high guarantees that all hypotheses receive gradients, but it may also overly homogenize the models and reduce diversity. A similar effect arises in the annealed variant, where the temperature scheduler parameters influence performance. Exploring dynamic adjustment of these parameter values depending on the data distribution is left for future work.
@@ Appendix A Notations and setup
In the following, $x\triangleq(x_{t})_{t=1}^{T}\in\mathcal{V}^{T}$ be a sequence of $T$ tokens belonging to a finite vocabulary $\mathcal{V}=\{1,\dots,|\mathcal{V}|\}$, and $c\triangleq(c_{t})_{t=1}^{\tau}\in\mathcal{C}$ be a sequence of $\tau$ context embeddings of dimension $d$. In the following $\mathcal{X}=\mathcal{V}^{T}$ and $\mathcal{C}=(\mathbb{R}^{d})^{\tau}$.
@@
In the following, we assume that the model $p_{\theta}$ is perfectly expressive. Formally, let $\mathcal{F}_{\Theta}\triangleq\{p_{\theta}:c\in\mathcal{C}\rightarrow p(\cdot\,|\,c)\in\mathcal{P}(\mathcal{X})\,|\,\theta\in\Theta\}$ be the family of conditional distributions realized by the model, where $\mathcal{P}(\mathcal{X})$ is the set of probability distributions on $\mathcal{X}$. We assume the family $\mathcal{F}_{\Theta}$ is perfectly expressive, that is $\mathcal{F}_{\Theta}=\mathcal{P}(\mathcal{X})^{\mathcal{C}}$.
First, note that we have the Proposition 2, a well-known result that is due to the Gibbs inequality (e.g., MacKay (2003)), for which we provide a proof for completeness.
Proposition 2 (MacKay, 2003).
@@
where $\mathcal{H}(x\,|\,c)\triangleq-\mathbb{E}_{c,x}[\mathrm{log}\;p(x\,|\,c)]$.
@@
$\int_{\mathscr{S}(p)}\mathrm{log}\left[\frac{p_{\theta}(x\,|\,c)}{p(x,c)}\right]p(x,c)\;\mathrm{d}x\mathrm{d}c\geq-\mathrm{log}\left(\int_{\mathscr{S}(p)}\frac{p_{\theta}(x\,|\,c)}{p(x,c)}p(x,c)\;\mathrm{d}x\mathrm{d}c\right)\geq 0\;.$ (12)
@@
$\int_{\mathcal{X}\times\mathcal{C}}p(x,c)\;\mathrm{log}\;p_{\theta}(x\,|\,c)\;\mathrm{d}x\mathrm{d}c\geq-\int_{\mathcal{X}\times\mathcal{C}}p(x,c)\;\mathrm{log}\;p(x\,|\,c)\;\mathrm{d}x\mathrm{d}c=\mathcal{H}(x\,|\,c)\;,$
@@
$\mathbb{E}_{\alpha}[\mathrm{log}\,\beta(x)]=\mathrm{KL}(\alpha\,\|\,\beta)+\mathcal{H}(\alpha)\;,$ (13)
@@ Appendix B Proof of Proposition 1
$D(\theta,q)\triangleq\int_{\mathcal{X}}\sum_{k=1}^{K}q(k\,|\,x)\mathrm{log}\,p(x,z_{k};\theta)\;\mathrm{d}p(x),\penalty 10000\ D(\theta)\triangleq\int_{\mathcal{X}}\max_{k=1,\dots,K}\;\mathrm{log}\,p(x,z_{k};\theta)\;\mathrm{d}p(x)\;,$ (16)
@@
Indeed, the expectation step comes down to computing $q(\theta^{(t)})$. For the Maximization step, we have: $D\left(\theta^{(t+1)},q(\theta^{(t)})\right)\geq D\left(\theta^{(t)},q(\theta^{(t)})\right)$, by definition. At the next expectation step, we have: $D\left(\theta^{(t+1)},q(\theta^{(t+1)})\right)\geq D\left(\theta^{(t+1)},q(\theta^{(t)})\right)$, because $q(\theta^{(t+1)})$ computes the best assignment given the parameters $\theta^{(t+1)}$. This shows that $D(\theta^{(t+1})\geq D(\theta^{(t)})$.
The main difference compared to the vanilla form of the (hard) EM algorithm is that $(i)$ the goal here is to fit a conditional distribution p(x|c)p(x\,|\,c) given pairs $(c,x)\sim p(c,x)$, and $(ii)$ step 2 performs a gradient update (of the neural network weights) instead of a full maximization.
@@
$\mathcal{L}^{\mathrm{WTA}}(\theta)=-\int_{\mathcal{X}\times\mathcal{C}}\max_{k=1,\dots,K}\mathrm{log}\;p(x\,|\,c;\theta_{k})\mathrm{d}p(c,x)\;.$ (17)
The loss is generally expected to decrease across training iterations, although strict monotonicity is not guaranteed without additional assumptions on the learning rate and the smoothness of the gradient of the loss (Bertsekas, 1997). Since the loss is bounded below (by 0), the sequence of loss values $\{\mathcal{L}^{\mathrm{WTA}}(\theta^{(t)})\}_{t\geq 0}$ is therefore expected to converge in practice.
To conclude, we can view $(\theta_{1},\dots,\theta_{K})$ as the parameters involved in the estimation of the modes of the conditional distribution with $p(x\,|\,c)=\sum_{k=1}^{K}p(x\,|\,c;\theta_{k})p(\theta_{k}\,|\,c)$. Note that the current form of the algorithm does not estimate the weight of each mode $p(\theta_{k}\,|\,c)$, further work could include incorporating scoring heads to estimate $p(\theta_{k}\,|\,c)$ each $k$ as in Letzelter et al. (2023). ∎
We then expect to be able to recover the distributions (with one permutation) $\{p(x\,|\,c,z_{k})\}$ from estimated $\{p(x\,|\,c;\theta_{k})\}$, assuming identifiability of the data generating mixture, which we expect to be made easier if the components are enough separated (e.g., Redner and Walker (1984) Par. 2.5 or McLachlan et al. (2019) Sec. 2.2).
Proof of (ii) Let us assume that (with one permutation) $p(x\,|\,z_{k},c)=p(x\,|\,c;\theta_{k})$ for $k\in\{1,\dots,K\}$. This is possible thanks to Assumption 1. Let us show that (14). Let us define:
@@
Now let us leverage Assumption 1 to choose $\tilde{\theta}_{k}$ such that $p(x\,|\,c,\tilde{\theta}_{k})=p(x\,|\,c,z_{k})$ for each $k\in\{1,\dots,K\}$. In this case, $\mathrm{max}_{k}\,p(x\,|\,c,\tilde{\theta}_{k})\geq p(x\,|\,c,z)$ for each $z\in\{z_{1},\dots,z_{K}\}$, and $\mathbb{E}_{z}\mathbb{E}_{x\,|\,z}\left[\mathrm{log}\frac{\mathrm{max}_{k}\,p(x\,|\,c,\tilde{\theta}_{k})}{p(x\,|\,c,z)}\right]\leq 0$. Then,